OpenStudy (anonymous):

Evaluate: -3 - (-2). I completely forgot how to do this but I do know PEMDAS is involved with solving the question. Can anyone refresh my memory for a medal?

OpenStudy (anonymous):

I don't need anyone too answer the question but tell me the first step.

OpenStudy (anonymous):

Are you adding or multiplying

OpenStudy (anonymous):

If you are adding or subtracting -(-2) would turn into +(+2)

OpenStudy (anonymous):

Thats the thing, I completely forgot. But I'm sure if someone tells me the first step I will remember.

OpenStudy (anonymous):

Which would be -3+2. If multiplying, distribute the - into -2 which will be +2. I guess both adding and subtracting have the same answer

OpenStudy (anonymous):

I do know how to negative times a negative equals a positive.

OpenStudy (anonymous):

-3+(-*-1)2=-3+2=-1

OpenStudy (anonymous):

I don't understand @zaibali.qasmi

OpenStudy (anonymous):

Oh I see it.

OpenStudy (anonymous):

But I just wanted to know the first step. I didn't want you to answer the question :P